OpenStudy (anonymous):

Write the standard equation of each Ellipse. Find the center, vertices, co-vertices and foci. Show all work 1. 25x^2 + 9y^2 = 225

OpenStudy (jessiegonzales):

Medal Pleas :) Simplifying 9y2 + -25x2 = 225 Reorder the terms: -25x2 + 9y2 = 225 Solving -25x2 + 9y2 = 225 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-9y2' to each side of the equation. -25x2 + 9y2 + -9y2 = 225 + -9y2 Combine like terms: 9y2 + -9y2 = 0 -25x2 + 0 = 225 + -9y2 -25x2 = 225 + -9y2 Divide each side by '-25'. x2 = -9 + 0.36y2 Simplifying x2 = -9 + 0.36y2 Reorder the terms: 9 + x2 + -0.36y2 = -9 + 0.36y2 + 9 + -0.36y2 Reorder the terms: 9 + x2 + -0.36y2 = -9 + 9 + 0.36y2 + -0.36y2 Combine like terms: -9 + 9 = 0 9 + x2 + -0.36y2 = 0 + 0.36y2 + -0.36y2 9 + x2 + -0.36y2 = 0.36y2 + -0.36y2 Combine like terms: 0.36y2 + -0.36y2 = 0.00 9 + x2 + -0.36y2 = 0.00

OpenStudy (jessiegonzales):

yes your welome?

OpenStudy (anonymous):

Thank you

OpenStudy (jessiegonzales):

medal?